# CHOPS: CHat with custOmer Profile Systems for Customer Service with LLMs

### v2

@@ Abstract
Businesses and software platforms are increasingly utilizing Large Language Models (LLMs) like GPT-3.5, GPT-4, GLM-3, and LLaMa-2 as chat assistants with file access or as reasoning agents for custom service. Current LLM-based customer service models exhibit limited integration with customer profiles and lack operational capabilities, while existing API integrations prioritize diversity over precision and error avoidance which are crucial in real-world scenarios for Customer Service. We propose an LLMs agent called CHOPS (CHat with custOmer Profile in existing System) that: (1) efficiently utilizes existing databases or systems to access user information or interact with these systems based on existing guidance; (2) provides accurate and reasonable responses or executing required operations in the system while avoiding harmful operations; and (3) leverages the combination of small and large LLMs together to provide satisfying performance while having decent inference cost. We introduce a practical dataset, CPHOS-dataset, including a database, some guiding files, and QA pairs collected from CPHOS, which employs an online platform to facilitate the organization of simulated Physics Olympiads for high school teachers and students. We conduct extensive experiments to validate the performance of our proposed CHOPS architecture using the CPHOS-dataset, aiming to demonstrate how LLMs can enhance or serve as alternatives to human customer service. Code for our proposed architecture and dataset can be found at https://github.com/JingzheShi/CHOPS.
@@ 1 Introduction
In most organizations with human customer service, a system usually stores customer information. Responses are based on the customer’s profile, like user type and purchase history, following set guidelines. Customer service can also update the customer’s status upon request. Large language models (LLMs), such as GPT-3.5 (OpenAI, 2022),GPT-4.0 (OpenAI, 2023), GLM (Zeng et al., 2022; Du et al., 2022)) and LLaMa(Touvron et al., 2023), have emerged as a representative achievement of AI development in the past decade. With their mastering of common knowledge and their ability to understand prompts and generate contextually relevant answers, LLMs have been used as assistants across a wide range of application scenarios, including chatting assistants, coding assistants, automatic assistant agents, etc(Kalla & Smith, 2023).
@@
A common paradigm for equipping LMs with external knowledge while avoiding long context length that are either costly or technically hard is RAG(Lewis et al., 2021), represented by the widely used and efficient Vector Database using a sentence embedding model such as the Universal Sentence Encoder(Cer et al., 2018). Most publicly available architectures for utilizing LLMs as customer service mainly follow this paradigm, e.g. Databricks(Databricks, ) enables users to upload guiding files thus building a Vector Database-based customer service agent to augment human customer service. However, for most software platforms or businesses, to answer based on a series of guides is not enough: to query information or to manipulate the system using a set of APIs is necessary in some scenarios.
@@
Figure 1: Left: Existing scenarios for Customer Service require File QA and System Manipulation. Middle: Possible mistakes in Customer Service. Accuracy is needed in this scenario, especially to avoid those harmful operations. Right: existing methods to use LLM as assistants. LLMs for APIs like ToolLLM (Qin et al., 2023) mainly focusing on a large number of APIs in API hubs.
@@
Previous works (involving models, agent architectures, and datasets) on LLMs using APIs (Patil et al., 2023; Qin et al., 2023; Tang et al., 2023) mainly focus on the LLM’s ability to choose between a vast amount (100 through 10000+) APIs and to accomplish different tasks. However, for a particular customer service scenario, much fewer APIs are needed but high accuracy are needed, especially for modifying user status in important aspects like banking. This difference in the focus point of Customer Service and previously defined API using tasks sheds light on a different method needed for Customer Service compared to previous API-using methods and new datasets for evaluation such methods.
@@
Online Customer Service, with its long history since the popularization of the Internet and personal computer, requires accurate answers or modifications to user profiles in customer service based on the user’s questions or requests and the guiding file. Unlike other scenarios for reasoning or prompting, especially on math problem solving  (Wei et al., 2023; Yao et al., 2023; Zhang et al., 2023), or utilizing numbers of APIs  (Patil et al., 2023; Qin et al., 2023; Tang et al., 2023), we focus on the new task of customer service, which sheds light on the accuracy and admittance of the answer; moreover, the cost of the LLM needs to be controlled.
@@
To this end, we propose a general framework called CHOPS: CHat with custOmer Profile in existing System as Figure 3 shown. In our CHOPS framework, we propose a classifier-executor-verifier based framework that make the LLM better utilize tools, especially Database and guidance. Through the work for these three agents, the task of answering to a user’s question is decomposed into (1) classifying the type or theme of the user’s question, (2) giving answers or deciding operations to be executed and (3) verify then reject or commit the result of the executor.
Our work makes several contributions to the application of LLMs, especially the field of doing customer service with LLMs.
@@ 2 Related Works
Incorporating external knowledge sources into LLMs for enhanced performance on knowledge-intensive tasks has seen advancements through Retrieval-Augmented Generation (RAG), with the use of vector-based databases for PDF files being a notable example Tripathi (2023). This approach encodes user queries into vectors, using k-nearest neighbors (KNN) to retrieve relevant information. In customer service, LLMs augmented with large databases have aimed to provide encyclopedic support for user inquiries dat (2022); Wulf (2022); cms (2022). However, such methods sometimes struggle in scenarios requiring modifications to a user’s profile within existing systems, a crucial aspect of customer service. This highlights the importance of integrating LLMs with software systems for direct interaction tasks, essential for operational efficiency in businesses.
@@
The research on LLMs as specialized agents is an evolving field in artificial intelligence and natural language processing. Initial research focused on using predefined prompts or fine-tuning to enhance LLMs for specific tasks, establishing their potential in specialized applications like natural language understanding. Recent studies (Hu et al., 2023a; Doe & Smith, 2023) explore LLMs in agent-based architectures for complex problem solving, such as mathematical puzzles, highlighting the importance of architectural design in improving performance. This progression suggests new avenues for leveraging LLMs in agent-based systems, offering insight into their capabilities for advanced interaction and problem solving tasks.
@@ 3.1 Database
The online system of the Simulated Olympiad utilizes a MySQL database. We provided 9 data desensitized table. The detailed description can be found in the appendix A.1.
In short, given a user’s nickname, one can do a series of queries on tables in the database to obtain or modify partially the profile of the user. The most important field of a user profile includes: (0) approved_to_use_online_platform; (1) user_name; (2) school_id; (3) user type: team leader, vise team leader, arbiter; (4) marking_question_id, etc.
Unlike previous works (Hu et al., 2023a) directly using LLMs to generate SQL commands for query or modification to the database, we wrapped the query and modification into a series of python APIs, following the idea of Repository Pattern in software design. We provide 9 Data Managing APIs and 18 Data Query APIs, 10 of which are available to LLMs. We collect instructions and queries to the system and augment them with GPT-4 into 104 System-related queries and instructions. There are several advantages of wrapping SQL commands and LLMs manipulate database through these APIs, in that:
  • Properly named APIs are much easier for LLMs to understand and to generate compared to complicated table structure and SQL commands for the database.
@@
  • In codes of software or websites that are written following the Repository Pattern (or more broadly, the principle of ‘encapsulation’ in software architecture), a series of pre-defined APIs for database is likely to exist. Much less effort is needed to modify these APIs into APIs suitable for LLMs than to check for correct SQL commands generated by LLMs.
@@ 3.2 PDF-based guides
There are two main guiding files provided by CPHOS: the mini-program guiding file and questions that are commonly asked by users, together with their answers. These QA files, together with the mini-program guiding file, is what we refer to as PDF-based guides. All files are translated by us into English. In practice we merge them into one file for RAG. Full pdf-based guides are appended in the supplementary materials. Collected QA pairs from real scene, we augment them through GPT-4 and repitition into 102 QA pairs on Guide Files, example of which can be shown in the left part of Figure 2.
@@ 4 Methods
Figure 3: Our CPHOS pipeline including Classifer, Executor and Verifier.
@@ 4.1 Framework Overview
We propose a three-agent architecture for the task: the classifier-executor-verifier architecture. For this three-stage architecture:
@@
    The Classifier is given the UserTexts, the System API descriptions and several relevant (and short) chunks from the guiding files. The Classifier classifies the UserTexts based on information needed for the following pipeline. The classifier itself does not output all relevant information, but only indicates what type of information is helpful.
@@
    The Executor is given the UserTexts and other information that the classifier classifies as helpful in the Classifier Stage. Note that the information can be richer than that given to the Classifier (e.g. longer retrieved chunks, more detailed descriptions, etc.). The Executor than gives a proposed answer or a proposed API call based on information given.
@@
    The Verifier is responsible of revisiting the Executor’s result and verify if it is valid or not and give a reason as well. If valid, then the answer is returned or the execution is carried out, and a reply is generated based on the result of that execution. If invalid, then the whole process will be redone, while the Classifier and the Executor can see the invalid reason provided by the verifier as a reference.
@@ 4.1.1 Input classifier
Previous work (Chen et al., 2023) has shown that the longer and more complicated the retrieved content is, the more difficult for the Executor LLM to find the exact piece of information and return it. Also, feeding the Executor everytime with retrieved chunks from guide file and API infos is token-consuming. We utilize a Classifier to classify the information domain that need to be given to the Executor in advance in order to solve this two issues.
A simple and experimentally-proved effective and efficient design for the Classifier is a binary classifier, or a 1-level Classifier as shown in the left part of Figure 4. This classifier only choose between two categories: (1) that the User Texts are about a query to the guiding file, and (2) that the User Texts are related to a query to the system or an instruction to the system. Non-classifiable cases are dealt with in the same way as (1) in the following pipeline. Note that although we use the same RAG method to retrieve file chunks for the Classifier and the Executor, the chunk length for Classifier is set to be lower so as to save inference cost.
Moreover, we observe that many queries and questions are about basic information of the user in the system that is much shorter and less token-consuming compared to the retrieved chunks. Inspired by the idea of Cache, we further add one categorized: (0) the ‘Basic Info’ apart from (1)‘Guide File’, (2)‘System API’. We design a 2-level Classifier Architecture shown in the right part of Figure 4. One classifier will first decide whether the User Texts are solvable only given the Basic Information (without the need of ‘Guide Files’ and ‘System APIs’). If it asserts yes, then the pipeline will go on and the Executor will only see the Basic Information. If not, then it goes to the second level classifier to categorize user texts into class (1) or class (2), and further provide the corresponding information to the Executor. This design also improves accuracy while saving token consumption.
@@ 4.1.3 Verifier
Previous works (Weng et al. (2023),Zhang et al. (2023)) have proved the effectiveness of a Verifier for re-checking correctness.
In our work, the Verifier verifies the result and, if valid, summarize the answer or generate a response to the user based on the executed operation.
@@
The Verifier is required to output valid score (1-10) and reasons at the same time. If the verification result is invalid, then we redo the whole C-E-V process while giving the Classifier and Executor with the invalid reason.
In practice to ensure fast response, we restrict the loop iterations into $5$. For latter iterations, a lower score would be seen as valid: we use a simple linear scheduling of the passing score with the iteration index. If all iterations fail, we ask the LLM to choose between one answers provided before.
@@ 4.2 Tools used
For PDF Retrieval, we follow the well-practised method of using a sentence encoder to encode chunks from guide files into a vector database and retrieve top-K closest chunks in latent space as related information given by the files. We follow pdfgpt(Tripathi, 2023) and utilize Universal Sentence Encoder (Cer et al., 2018) as the embedding model. For Database manipulation, we use the wrapped Apis as available tools.
@@ 5.2 Main Experiment Results
C-E-V gpt-3.5-turbo 95.2 90.2 69.5% 30.1 0.56
C-E-V Mixed 98.0 99.0 116.4% 16.86+9.79 0.33+0.21
Table 1: Main Experiment Result.: gpt-4-0125-preview is used. : gpt-3.5-turbo is used for Classifier and Verifier, while gpt-4-0125-preview is used for executor. gpt-3-turbo characters + gpt-4-0125-preview characters. Pricing for calculating relative cost is the price obtained from OpenAI at Mar.2024. See Appendix A.2 for more details about calculating relative cost.
We evaluate the proposed CHOPS agent architecture on the CPHOS-dataset. We conduct prompt-tuning on gpt-4 (gpt-4-0125-preview) as our baseline, which is labeled as Executor Only in our result tables. We show our main experiment results in Table 1.
In all, our proposed C-E-V agent architecture with gpt-3.5-turbo as LLM backbones for all agents surpasses this gpt-4 baseline by a large margin ($86\%\rightarrow 95\%,83\%\rightarrow 90\%$) with $69.5\%$ cost. Moreover, by substituting the Executor backbone with gpt-4, we can reach accuracy of above $98\%$ on both accuracy metrics, hugely surpassing the plain gpt-4 baseline while costing only $16\%$ more than the plain gpt-4 solution (see Appendix A.2 for more details about price estimation). We claim our architecture is flexible and is token-efficient, reaching a balance between cost and accuracy.

### v4

@@ Abstract
Businesses and software platforms are increasingly using large language models (LLMs) such as GPT-3.5, GPT-4, GLM-3, and LLaMa-2 as chat assistants with file access or as reasoning agents for custom service. Current LLM-based customer service models exhibit limited integration with customer profiles and lack operational capabilities, while existing API integrations prioritize diversity over precision and error avoidance, which are crucial in real-world scenarios for Customer Service. We propose an LLMs agent called CHOPS (CHat with custOmer Profile in existing System) that: (1) efficiently utilizes existing databases or systems to access user information or interact with these systems based on existing guidance; (2) provides accurate and reasonable responses or executing required operations in the system while avoiding harmful operations; and (3) leverages the combination of small and large LLMs together to provide satisfying performance while having decent inference cost. We introduce a practical dataset, CPHOS-dataset, including a database, some guiding files, and QA pairs collected from CPHOS*. We conduct extensive experiments to validate the performance of our proposed CHOPS architecture using the CPHOS-dataset, aiming to demonstrate how LLMs can enhance or serve as alternatives to human customer service. Code for our proposed architecture and dataset can be found at https://github.com/JingzheShi/CHOPS.
@@ 1 Introduction
In most organizations with human customer service, there is usually a system storing customer information. Responses are based on the customer’s profile, like user type and purchase history, following set guidelines. Customer service can also update the customer’s status upon request. Large language models (LLMs), such as GPT-3.5 (OpenAI, 2022), GPT-4.0 (OpenAI, 2023), GLM (Zeng et al., 2022; Du et al., 2022) and LLaMa (Touvron et al., 2023), have emerged as a representative achievement of AI development in the past decade. With their mastering of common knowledge and their ability to understand prompts and generate contextually relevant answers, LLMs have been used as assistants across a wide range of application scenarios, including chatting assistants, coding assistants, automatic assistant agents, etc (Kalla & Smith, 2023).
@@
A common paradigm for equipping LMs with external knowledge while avoiding long context lengths that are costly or technically hard is RAG (Lewis et al., 2021), represented by the widely used and efficient Vector Database using a sentence embedding model such as the Universal Sentence Encoder (Cer et al., 2018). Most publicly available architectures for utilizing LLMs as customer service mainly follow this paradigm, e.g. Databricks (Databricks, ) enables users to upload guiding files thus building a Vector Database-based customer service agent to augment human customer service. However, for most software platforms or businesses, to answer based on a series of guides is not enough: to query information or to manipulate the system using a set of APIs is necessary in some scenarios.
@@
Figure 1: Left: Existing scenarios for Customer Service require File QA and System Manipulation. Middle: Possible mistakes in Customer Service. Accuracy is needed in this scenario, especially to avoid those harmful operations. Right: existing methods to use LLM as assistants. LLMs for APIs like ToolLLM (Qin et al., 2023) mainly focus on a large number of APIs in API hubs.
@@
Previous works (involving models, agent architectures, and datasets) on LLMs using APIs (Patil et al., 2023; Qin et al., 2023; Tang et al., 2023) mainly focus on the LLM’s ability to choose between a vast amount (100 through 100,000+) APIs and to accomplish different tasks. However, for a particular customer service scenario, much fewer APIs are needed but high accuracy is needed, especially for modifying user status in important aspects like banking. This difference in the focus point of Customer Service and previously defined API-using tasks sheds light on a different method needed for Customer Service compared to previous API-using methods and new datasets for evaluation of such methods.
@@
Online Customer Service, with its long history since the popularization of the Internet and personal computers, requires accurate answers or modifications to user profiles in customer service based on the user’s questions or requests and the guiding file. Unlike other scenarios for reasoning or prompting, especially on math problem solving (Wei et al., 2023; Yao et al., 2023; Zhang et al., 2023), or utilizing numbers of APIs (Patil et al., 2023; Qin et al., 2023; Tang et al., 2023), we focus on the new task of customer service, which sheds light on the accuracy and admittance of the answer; moreover, the cost of the LLM needs to be controlled.
@@
To this end, we propose a general framework called CHOPS: CHat with custOmer Profile in existing System as Figure 3 shown. In our CHOPS framework, we propose a classifier-executor-verifier based framework that makes the LLM better utilize tools, especially Database and guidance. Through the work of these three agents, the task of answering a user’s question is decomposed into (1) classifying the type or theme of the user’s question, (2) giving answers or decision operations to be executed and (3) verifying then reject or commit the result of the executor.
Our work makes several contributions to the application of LLMs, especially in the field of customer service with LLMs.
@@ 2 Related Works
Incorporating external knowledge sources into LLMs for enhanced performance on knowledge-intensive tasks has seen advancements through Retrieval-Augmented Generation (RAG), with the use of vector-based databases for PDF files being a notable example (Tripathi, 2023). This approach encodes user queries into vectors, using k-nearest neighbors (KNN) to retrieve relevant information. In customer service, LLMs augmented with large databases have aimed to provide encyclopedic support for user inquiries (dat, 2022; Wulf, 2022; cms, 2022). However, such methods sometimes struggle in scenarios requiring modifications to a user’s profile within existing systems, a crucial aspect of customer service. This highlights the importance of integrating LLMs with software systems for direct interaction tasks, essential for operational efficiency in businesses.
@@
The research on LLMs as specialized agents is an evolving field in artificial intelligence and natural language processing. Initial research focused on using predefined prompts or fine-tuning to enhance LLMs for specific tasks, establishing their potential in specialized applications like natural language understanding. Recent studies (Hu et al., 2023a; Doe & Smith, 2023) explore LLMs in agent-based architectures for complex problem solving, such as mathematical puzzles, highlighting the importance of architectural design in improving performance. This progression suggests new avenues for leveraging LLMs in agent-based systems, offering insight into their capabilities for advanced interaction and problem-solving tasks.
@@ 3.1 Database
The online system of the Simulated Olympiad utilizes a MySQL database. We provided 9 data desensitized tables. The detailed description can be found in the appendix A.1.
In short, given a user’s nickname, one can do a series of queries on tables in the database to obtain or modify partially the profile of the user. The most important field of a user profile includes (0) approved_to_use_online_platform; (1) user_name; (2) school_id; (3) user type: team leader, vise team leader, arbiter; (4) marking_question_id, etc.
Unlike previous works (Hu et al., 2023a) directly using LLMs to generate SQL commands for query or modification to the database, we wrapped the query and modification into a series of python APIs, following the idea of Repository Pattern in software design. We provide 9 Data Managing APIs and 18 Data Query APIs, 10 of which are available to LLMs. We collect instructions and queries to the system and augment them with GPT-4 into 104 System-related queries and instructions. There are several advantages of wrapping SQL commands and LLMs manipulate databases through these APIs, in that:
  • Properly named APIs are much easier for LLMs to understand and to generate compared to complicated table structures and SQL commands for the database.
@@
  • In codes of software or websites that are written following the Repository Pattern (or more broadly, the principle of ‘encapsulation’ in software architecture), a series of pre-defined APIs for a database is likely to exist. Much less effort is needed to modify these APIs into APIs suitable for LLMs than to check for correct SQL commands generated by LLMs.
@@ 3.2 PDF-based guides
There are two main guiding files provided by CPHOS: the mini-program guiding file and questions that are commonly asked by users, together with their answers. These QA files, together with the mini-program guiding file, are what we refer to as PDF-based guides. All files are translated by us into English. In practice, we merge them into one file for RAG. Full PDF-based guides are appended in the supplementary materials. Collected QA pairs from real scene, we augment them through GPT-4 and repetition into 102 QA pairs on Guide Files, an example of which can be shown in the left part of Figure 2.
@@ 4 Methods
Figure 3: Our CHOPS architecture including Classifer, Executor and Verifier.
@@ 4.1 Framework Overview
We propose a three-agent architecture for this task, termed the classifier-executor-verifier architecture. In this architecture, each agent is implemented using a Large Language Model (LLM). This three-stage architecture functions as follows:
@@
    The Classifier is given the UserTexts, the System API descriptions and several relevant (and short) chunks from the guiding files. The Classifier classifies the UserTexts based on information needed for the following pipeline. The classifier itself does not output all relevant information but only indicates the type of information that would be beneficial for further processing.
@@
    The Executor is given the UserTexts and other information that the classifier classifies as helpful in the Classifier Stage. Note that the information can be richer than that given to the Classifier (e.g. longer retrieved chunks, more detailed descriptions, etc.). The Executor then gives a proposed answer or API call based on the information given.
@@
    The Verifier is responsible for revisiting the Executor’s result, verifying if it is valid or not and giving a reason as well. If valid, then the answer is returned or the execution is carried out, and a reply is generated based on the result of that execution. If invalid, then the whole process will be redone, while the Classifier and Executor can see the invalid reason provided by the verifier as a reference.
@@ 4.1.1 Input Classifier
Previous work (Chen et al., 2023) has shown that the longer and more complicated the retrieved content, the more difficult it is for the Executor LLM to find the exact piece of information and return it. Also, feeding the Executor every time with retrieved chunks from guide file and API info is token-consuming. We utilize a Classifier to classify the information domain that needs to be given to the Executor in advance to solve these two issues.
A simple and experimentally-proved effective and efficient design for the Classifier is a binary classifier, or a 1-level Classifier as shown in the left part of Figure 4. This classifier only chooses between two categories: (1) that the UserTexts are about a query to the guiding file, and (2) that the UserTexts are related to a query or an instruction to the system. Non-classifiable cases, which are not related to the above two cases, are dealt with in the same way as (1) in the following pipeline. Note that although we use the same RAG method to retrieve file chunks for the Classifier and the Executor, the chunk length for Classifier is set to be lower to save inference cost.
Moreover, we observe that many queries and questions are about basic information of the user in the system that is much shorter and less token-consuming compared to the retrieved chunks. Inspired by the idea of Cache, we further add one categorized: (0) ‘Basic Info’ apart from (1) ‘Guide File’, (2) ‘System API’. This leads to the development of a two-level classifier architecture shown in the right part of Figure 4. One classifier will first decide whether the UserTexts are solvable only given the Basic Information (without the need for ‘Guide Files’ and ‘System APIs’). If it asserts yes, then the pipeline will go on and the Executor will only see the Basic Information. If not, then it goes to the second level classifier to categorize UserTexts into class (1) or (2), and further provide the corresponding information to the Executor. The experiment indicates this architecture improves accuracy while saving token consumption.
@@ 4.1.3 Verifier
Previous works (Weng et al., 2023; Zhang et al., 2023) have proved the effectiveness of a Verifier for re-checking correctness.
In our work, the Verifier verifies the result and, if valid, summarizes the answer or generates a response to the user based on the executed operation.
@@
The Verifier is required to output a valid score (1-10) and reasons at the same time. If the verification result is invalid, then we redo the whole C-E-V process while giving the Classifier and Executor the invalid reason.
In practice to ensure fast response, we restrict the loop iterations into $5$. For latter iterations, a lower score would be seen as valid: we use a simple linear scheduling of the passing score with the iteration index. If all iterations fail, we ask the LLM to choose between one answer provided before. More details about failure cases and number of iterations in practice can be found in Appendix A.4.
@@ 4.2 Tools used
For PDF Retrieval, we follow the well-practiced method of using a sentence encoder to encode chunks from guide files into a vector database and retrieve top-K closest chunks in latent space as related information given by the files. We follow Pdf-gpt (Tripathi, 2023) and utilize Universal Sentence Encoder (Cer et al., 2018) as the embedding model. For Database manipulation, we use the wrapped APIs as available tools.
@@ 5.1 Metrics
We make a comparison at the character level since different LLMs utilize different tokenization methods. Input characters and output characters are separated since generating output tokens is more resource-consuming than reading input tokens for LLMs (and are more expensive in terms of API price). For clarity, we define the first two metrics mathematically.
$\text{Accuracy}=\frac{N_{correct}}{N_{total}}$ (1)
Where:
  • $N_{correct}$ is the number of questions correctly answered by the model.
  • $N_{total}$ is the total number of questions and instructions.
@@ 5.2 Main Experiment Results
1-vote CoT gpt-3.5-turbo 91.3 62.7 16.9% 14.51 0.5
4-vote CoT gpt-3.5-turbo 90.2 77.4 61.0% 53.54 0.94
16-vote CoT gpt-3.5-turbo 90.2 78.4 239.5% 211.58 2.59
C-E-V gpt-3.5-turbo 95.2 90.2 34.4% 30.1 0.56
C-E-V Mixed 98.0 99.0 96.6% 16.86+9.79 0.33+0.21
Table 1: Main Experiment Result.: gpt-4-0125-preview is used. : gpt-3.5-turbo is used for Classifier and Verifier, while gpt-4-0125-preview is used for Executor. : gpt-3-turbo characters + gpt-4-0125-preview characters. Pricing for calculating relative cost is the price obtained from OpenAI in Mar.2024. See Appendix A.2 for more details about calculating relative cost.
We evaluate the proposed CHOPS agent architecture on the CPHOS-dataset. We perform prompt tuning on gpt-4 (gpt-4-0125-preview, labeled Executor Only) and use multivote CoT (on gpt-3.5-turbo) as our baselines. We show our main experiment results in Table 1.
Compared to the CoT method, our method achieves higher accuracy while maintaining a decent cost. Multi-vote CoT does see an improvement in performance when increasing the number of votes, but it cannot provide satisfactory performance even with more cost.
Moreover, our proposed C-E-V agent architecture with gpt-3.5-turbo as LLM backbones for all agents surpasses the Executor Only gpt-4 baseline by a large margin ($86\%\rightarrow 95\%,83\%\rightarrow 90\%$) with $34.4\%$ cost. Moreover, by substituting the Executor backbone with gpt-4, we can reach accuracy of above $98\%$ on both accuracy metrics, hugely surpassing the plain gpt-4 baseline while even cost less than the gpt-4 baseline (see Appendix A.2 for more details about price estimation). Thus our architecture is proven to be flexible and is token-efficient, reaching a balance between cost and accuracy.